# all_papers_200910.bib

### v2

@@ Abstract
Corresponding author: Shenhao Wang; email: shenhaowang@ufl.eduPaper is published in Transportation Research Part B: Methodological with the link as https://www.sciencedirect.com/science/article/pii/S0191261524001851
@@ 1 Introduction
In recent years, there has been a growing interest in incorporating machine learning (ML) models into travel demand analysis, alongside the traditional choice models (DCMs) that have been widely used for decades [Ben_Akiva1996, Cantarella2005, Small2007td, Ortuzar2011]. These studies seek to compare the differences in model performance, system robustness, and ease of interpretation between ML and DCMs [Hagenauer2017, WangFangru2018, ChengLong2019]. These comparative studies play a crucial role in generating valuable insights into novel behavioral mechanisms and diversifying travel demand models in practical applications [WangShenhao2020_econ_info, WangShenhao2020_asu_dnn].
@@
To address the limitations, we develop a large-scale empirical benchmark by designing a tournament model to compare the performance of hundreds of ML and DCMs. The tournament model uses the winning model in an experiment as the output, and alternative models and contextual factors as the inputs [^1] Inputs and outputs are commonly used in the ML literature, and they correspond to independent variables and dependent variables in choice modeling.. It captures the essence of comparative experiments by conducting pairwise comparison for the intrinsic predictive values of models using statistical methods. With a formal statistical approach, our benchmark study can efficiently summarize findings, differentiate between the model and contextual effects, and highlight the shared patterns from a large number of comparative experiments. With the tournament model, this benchmark study can filter out the context-specific noises and reveal common patterns in a specific topic, resembling a conventional meta-analysis while strengthening it with our own large-scale experiments [Ewing2010, Holmgren2007, Patricia2009]. As far as we know, such an empirical benchmark is still missing for travel demand modeling.
@@ 3 Literature data
1 Nijkamp et al. (1996) [Nijkamp1996] MC $1,396$ 2 - DNN, MNL DNN
2 Rao et al. (1998) [Rao1998] MC $4,335$ 2 - DNN, MNL DNN
3 Hensher and Ton (2000) [Hensher2000] MC $1,500$ 2 - DNN, NL DNN/NL
4 Sayed and Razavi (2000) [Sayed2000] MC $7,500$ 2 - MNL, DNN DNN
5 Cantarella et al. (2002) [Cantarella2002] MC $2,808$ 3 - DNN, MNL, NL DNN
6 Mohammadian et al. (2002) [Mohammadian2002] CO $597$ 2 - DNN, NL DNN
7 Doherty and Mohammadian (2003) [Doherty2003] TP $5,583$ 2 - DNN, GLM DNN
8 Xie et al. (2003) [XieChi2003] MC $,74$ 747 3 - DT, DNN, MNL DNN
9 Cantarella et al. (2005) [Cantarella2005] MC $1,067$ 2 - DNN, MNL DNN
10 Celikoglu (2006) [Celikoglu2006] MC $381$ 4 - DNN, MNL DNN
11 Tortum et al. [Tortum2009] MC $441$ 3 - MNL, GLM, DNN DNN
12 Zhang and Xie (2008) [ZhangYunlong2008] MC $1,000$ 3 - MNL, DNN, SVM SVM
12 Zhang and Xie (2008) [ZhangYunlong2008] MC $2,000$ 3 - MNL, DNN, SVM SVM
13 Biagioni et al. (2008) [Biagioni2008] MC $19,118$ 4 - MNL, BM, BOOSTING, DT BOOSTING
14 Xian and Jian (2011) [XianYu2011] MC $4,725$ 3 - SVM, NL, DNN SVM
15 Allahviranloo and Recker (2013) [Allahviranloo2013] TP $3,671$ 2 - SVM, MNL SVM
16 Omrani et al. (2013) [Omrani2013] MC $3,673$ 6 - DNN, SVM, KNN, MNL, DT, BM DNN
17 Ermagun et al. (2015) [Ermagun2015] MC $4,700$ 2 - NL, RF RF
18 Jahangiri et al. (2015) [Jahangiri2015] MC N.A. 6 - KNN, SVM, DT, BAGGING, RF, MNL RF
19 Tang et al. (2015) [Tang2015] MC $72,536$ 2 - DT, MNL DT
20 Omrani (2015) [Omrani2015] MC $9,500$ 4 - DNN, RBFNN, MNL, SVM DNN
21 Shafique et al. (2015) [Shafique2015] MC $1,968$ 4 - SVM, BOOSTING, DT, RF RF
21 Shafique et al. (2015) [Shafique2015] MC $1,488$ 4 - SVM, BOOSTING, DT, RF RF
21 Shafique et al. (2015) [Shafique2015] MC $2,754$ 4 - SVM, BOOSTING, DT, RF RF
22 Shukla et al. (2015) [Shukla2015] MC $100,000$ 2 - DNN, DT DNN
23 Sekhar and Madhu (2016) [Sekhar2016] MC $4,976$ 3 - RF, DT, MNL RF
24 Hagenauer and Helbich (2017) [Hagenauer2017] MC $230,608$ 8 - MNL, DNN, NB, SVM, CTs, BOOSTING, BAGGING, RF RF
25 Paredes et al. (2017) [Paredes2017] CO $15,211$ 5 - MNL, BOOSTING, DT, SVM, RF RF
26 Hillel et al. (2018) [Hillel2018] MC N.A. 8 - DNN, BAGGING, BOOSTING, KNN, GLM, BM, RF, SVM BOOSTING
27 Golshani et al. (2018) [Golshani2018] MC $9,450$ 2 - MNL, DNN DNN
@@ 23 Sekhar and Madhu (2016) [Sekhar2016] MC $4,976$ 3 - RF, DT, MNL RF
29 Wang and Ross (2018) [WangFangru2018] MC $51,910$ 2 - BOOSTING, MNL BOOSTING
@@
31 Cheng et al. (2019) [ChengLong2019] MC $7,276$ 4 - RF, SVM, BOOSTING, MNL RF
32 Zhou et al. (2019) [ZhouXiaolu2019] MC $30,000$ 8 - MNL, KNN, DT, SVM, BM, BOOSTING, BAGGING, RF BAGGING
33 Wang et al. (2020) [WangShenhao2020_asu_dnn] MC $8,418$ 8 - MNL, NL, DNN, SVM, BM, KNN, BOOSTING, GLM DNN
33 Wang et al. (2020) [WangShenhao2020_asu_dnn] MC $,92$ 929 8 - MNL, NL, DNN, SVM, BM, KNN, BOOSTING, GLM DNN
34 Wang et al. (2020) [WangShenhao2020_econ_info] MC $80,000$ 2 - MNL, DNN DNN
34 Wang et al. (2020) [WangShenhao2020_econ_info] MC $8,418$ 2 - MNL, DNN DNN
35 Wang et al. (2020) [WangShenhao2020_rpsp] MC $8,418$ 3 - NL, MNL, DNN DNN
@@ 4.1 Treatment dimension: models and model families
Table LABEL:table:list_classifiers summarizes the list of 105 individual models (Column 1) from 12 model families (Column 2). Instead of analyzing single models, we create a two-tier structure for models and model families, thus recognizing the richness within each model family. For example, the DNN family includes 16 DNNs with different architectures, thus enabling us to analyze the variation across the DNN architectures. While far from exhaustive, Table LABEL:table:list_classifiers has included the relatively representative models and the most relevant ones for travel behavioral analysis. For example, DCMs are highly relevant given their consistent use in the travel demand analysis, and DNNs are also highly relevant due to their rising popularity in many subdomains in transportation [Karlaftis2011, WangShenhao2020_asu_dnn, WangShenhao2020_rpsp].
@@ 2. Deep Neural Networks (15 Models)
avNNet_R DNN Neural network with random seeds with averaged scores; [Ripley1996] R Caret avNNet
@@
monmlp_R DNN MLP with monotone constraints [ZhangHong1999] R Caret monmlp
@@ 3. Discriminant Analysis (12 Models)
pda_R DA Penalized discriminant analysis (PDA) with shrinkage penalty coefficients [Hastie1995] R mda gen.ridge
mda_R DA Mixture discriminant analysis (MDA) where the number subclass is tuned to 3 [Hastie1996] R mda
rda_R DA Regularized discriminant analysis (RDA) with regularized group covariance matrices [Friedman1989] R klaR
hdda_R DA High dimensional discriminant analysis (hdda) assuming each class in a Gaussian subspace [Bouveyron2007] R HD
@@ 4. Bayesian Models (7 Models)
BayesNet_W BM Bayes network models by hill climbing algorithm [Cooper1992] Weka BayesNet
@@ 6. K Nearest Neighbors (4 Models)
lBk_1_W KNN KNN classifier with number of neighbors equal to 1 (brute force searching and Euclidean distance) [Aha1991] Weka lBk
lBk_5_W KNN KNN classifier with number of neighbors equal to 5 (brute force searching and Euclidean distance) [Aha1991] Weka lBk
@@ 7. Decision Tree (14 Models)
C5.0Rules_R DT Rule-based models using Quinlan’s C5.0 algorithm [quinlan1993c4] R Caret C5.0Rules
ctree_R DT Conditional inference trees [Hothorn2006] R Caret ctree
@@
ExtraTree_P DT Tree classifier with best splits and features chosen from random splits and randomly selected features [Geurts2006] Python sklearn ExtraTreeClassifier
@@
HoeffdingTree_W DT An incremental tree with inductive algorithm. [Hulten2001] Weka HoeffdingTree
@@
DecisionTable_W DT Simple decision table majority classier that uses BestFirst as search method [Kohavi1995] Weka DecisionTable
@@ 8. Generalized Linear Models (10 Models)
Logistic_W GLM Logistic regression model with a ridge estimator [LeCessie1992] Weka Logistic
SimpleLogistic_W GLM Linear logistic regression models fitted by using LogitBoost [Landwehr2005] Weka SimpleLogistic
@@
Passive Aggressive_P GLM Passive-aggressive algorithms for classification with hinge loss [Crammer2006] Python sklearn PassiveAggressiveClassifier
@@ 12. Boosting (23 Models)
AdaBoost_P BOOSTING AdaBoost classifier. The DecisionTree with maximum depth =10 is set as the base estimator. [Freund1997] Python sklearn AdaBoostClassifier
@@
Gradient Boosting_P BOOSTING An additive model trained in a forward stage-wise fashion [Friedman2001] Python sklearn GradientBoostingClassifier
@@ 4.2 Contextual factors
The contextual variables are data sets, sample sizes, and choice categories, which can also influence the model performance besides the model choice. The data sets include NHTS2017, LTDS2015, and SGP2017. NHTS2017 refers to the national household travel survey 2017, collected by the Federal Highway Administration [^3] Available at https://nhts.ornl.gov/, which provides travel, vehicle, and individual information across the United States. LTDS2015 refers to the London travel demand survey 2015, which was collected in London with the history of trips from April 2012 to March 2015 and augmented with mode-specific level-of-service variables (e.g. in-vehicle travel time, public transport fares, fuel cost, etc.) [hillel2018recreating]. SGP2017 was collected in 2017 in Singapore through a stated preference survey for travel mode choice [WangShenhao2019_risk_av, shen2019built, mo2020Impacts], in which the experiments followed the orthogonal survey design [louviere2000stated]. SGP2017 contains information about travel mode choices, socio-demographics, and alternative specific variables. The three data sets are chosen to cover a variety of data collection procedures, geographical locations, and local contexts. The NHTS2017 was collected through revealed preference surveys by the US government, LTDS2015 by combining a transit agency’s survey with simulation-based travel information, and SGP2017 by a standard stated preference survey. The three datasets cover the continents of America, Europe, and Asia, thus jointly creating a geographically diverse set.
@@
The choice categories include three travel behavior variables: travel modes, trip purposes, and car ownership, which are amongst the most common choice categories from the literature. Travel mode choice is available for all three data sets, while trip purposes and car ownership are only available for the NHTS2017 data. In the NHTS2017 data, the initial 21 travel mode categories are aggregated into six major modes: walk and bike, car, SUV, van and track, public transit, and others. The NHTS2017 data has five trip purposes, including home-based work, home-based shopping, home-based social, home-based others, and none-home-based trips. The car ownership in the NHTS2017 data also has five categories, ranging from zero to more than three cars. The outputs from the three data sets are summarized in Figure 4.2. In the NHTS2017 data, the most common travel mode is car, accounting for 45.3%, while only 1.8% of the trips are taken by public transit. In the LTDS2015 and SGP2017 data sets, the largest mode share is driving, which accounts for 44.1% and 44.7%, respectively. Public transit is the second largest mode share in London and Singapore, accounting for 35.3% and 23.0% of the total trips.
@@
(a) NHTS2017-MC
(b) NHTS2017-TP
(c) NHTS2017-CO
(d) LTDS2015-MC
(e) SGP2017-MC
Fig. 1.: Output distributions in three datasets
@@ 5.2 Examining performance across model families and individual models
Figures 5.2 and 5.2 present visualizations of prediction accuracy for 12 model families and 105 individual models. Both figures are sorted based on average prediction accuracy, from highest to lowest. In Figure 5.2, the x-axis represents model families, while the y-axis represents prediction accuracy. Each white dot represents an experiment’s accuracy, and the grey areas indicate density distributions of predictive performance. The blue, red, and green dots represent the mean, median, and maximum prediction accuracy, respectively. Figure 5.2 extends the 12 model families to 105 classifiers using a similar format, replacing the violin plots with simple blue bars to depict performance variance. In this figure, the blue and red dots respectively represent the mean and median prediction accuracy. Additionally, Appendix IV contains an analysis of literature data, yielding similar findings to the experiment data presented here.
Fig. 2.: Prediction accuracy of 12 model families
@@
Figure 5.2 illustrates the relative prediction accuracy of different model families. Among them, ensemble methods and DNNs exhibit relatively high prediction accuracy, while the DCM family ranks 7th out of 12. The DNN family, comprising 15 classifiers, achieves mean and median accuracies of 57.79% and 56.77% respectively. The three ensemble methods, namely random forests, boosting, and bagging, also demonstrate high accuracy with mean values of 57.05%, 57.03%, and 55.42% respectively. Despite incorporating more complex error structures in the NL and MXL models, the DCMs still fail to outperform the ensemble methods and DNNs. The superior prediction accuracy of ML models can be attributed to their strong approximation power [Hornik1989] and effective regularization techniques [Srivastava_Hinton2014, Boyd2004]. This visualization further confirms the findings from our previous tournament model, providing additional ranking information for each model family.
Fig. 3.: Prediction accuracy of 105 models
@@
Figure 5.2 also reveals significant variation and even multi-modal distributions in predictive performance. For instance, the DNN family exhibits an average accuracy of approximately 57.79%, while its maximum prediction accuracy reaches around 74.61%. Similarly, the average accuracy of the DCM family stands at 54.43%, which is 15.37% lower than its maximum value of 69.8%. Remarkably, each model family can achieve maximum values in the range of 70-80%, surpassing the average by 20-30 percentage points. Several model families display explicit multi-modal distributions, including DCM and BAGGING. The wide variation and multi-modal distributions in prediction accuracy visually demonstrate that the choice of models may not be the most crucial factor in determining the final predictive performance. This observation aligns with our previous finding regarding the substantial uncertainty inherent in such model comparisons.
@@
Figure 5.2 demonstrates that individual models do not deviate significantly from the average performance of their respective model families, and substantial variation persists within each of the 105 individual models. Among these models, the top performers are LogitBoosting and GradientBoosting from the BOOSTING family, followed by DNNs such as nnet, avNNet, and monmlp, which rank 4th, 5th, and 10th out of 105 models. Within the DCM model family, MNL, NL, and MXL (highlighted in red) exhibit similar performance at the medium to lower end of the spectrum among the 105 models. Their average prediction accuracies are 54.50%, 55.01%, and 52.58%, respectively, only slightly deviating from the DCM family’s average. However, each individual model represents a significant variation, with performance varying by at least 20 percentage points, as indicated by the long blue bars. The fact that MXL performs worst may come as a surprise to readers, given that in estimation, the inclusion of random heterogeneity almost invariably leads to big gains in performance. However, here, we are talking about prediction performance, which implies averaging again over the random components, and this removes much of the differences between MXL and non-mixture DCM models.

### v3

# Comparing hundreds of machine learning and discrete choice models for travel demand modeling: an empirical benchmark

@@ 1 Introduction
In recent years, there has been a growing interest in incorporating machine learning (ML) models into travel demand analysis, alongside the traditional choice models (DCMs) that have been widely used for decades [1, 2, 3, 4]. These studies seek to compare the differences in model performance, system robustness, and ease of interpretation between ML and DCMs [5, 6, 7]. These comparative studies play a crucial role in generating valuable insights into novel behavioral mechanisms and diversifying travel demand models in practical applications [8, 9].
@@
To address the limitations, we develop a large-scale empirical benchmark by designing a tournament model to compare the performance of hundreds of ML and DCMs. The tournament model uses the winning model in an experiment as the output, and alternative models and contextual factors as the inputs [^1] Inputs and outputs are commonly used in the ML literature, and they correspond to independent variables and dependent variables in choice modeling.. It captures the essence of comparative experiments by conducting pairwise comparison for the intrinsic predictive values of models using statistical methods. With a formal statistical approach, our benchmark study can efficiently summarize findings, differentiate between the model and contextual effects, and highlight the shared patterns from a large number of comparative experiments. With the tournament model, this benchmark study can filter out the context-specific noises and reveal common patterns in a specific topic, resembling a conventional meta-analysis while strengthening it with our own large-scale experiments [10, 11, 12]. As far as we know, such an empirical benchmark is still missing for travel demand modeling.
@@ 3 Literature data
1 Nijkamp et al. (1996) [13] MC $1,396$ 2 - DNN, MNL DNN
2 Rao et al. (1998) [14] MC $4,335$ 2 - DNN, MNL DNN
3 Hensher and Ton (2000) [15] MC $1,500$ 2 - DNN, NL DNN/NL
4 Sayed and Razavi (2000) [16] MC $7,500$ 2 - MNL, DNN DNN
5 Cantarella et al. (2002) [17] MC $2,808$ 3 - DNN, MNL, NL DNN
6 Mohammadian et al. (2002) [18] CO $597$ 2 - DNN, NL DNN
7 Doherty and Mohammadian (2003) [19] TP $5,583$ 2 - DNN, GLM DNN
8 Xie et al. (2003) [20] MC $,74$ 747 3 - DT, DNN, MNL DNN
9 Cantarella et al. (2005) [2] MC $1,067$ 2 - DNN, MNL DNN
10 Celikoglu (2006) [21] MC $381$ 4 - DNN, MNL DNN
11 Tortum et al. [22] MC $441$ 3 - MNL, GLM, DNN DNN
12 Zhang and Xie (2008) [23] MC $1,000$ 3 - MNL, DNN, SVM SVM
12 Zhang and Xie (2008) [23] MC $2,000$ 3 - MNL, DNN, SVM SVM
13 Biagioni et al. (2008) [24] MC $19,118$ 4 - MNL, BM, BOOSTING, DT BOOSTING
14 Xian and Jian (2011) [25] MC $4,725$ 3 - SVM, NL, DNN SVM
15 Allahviranloo and Recker (2013) [26] TP $3,671$ 2 - SVM, MNL SVM
16 Omrani et al. (2013) [27] MC $3,673$ 6 - DNN, SVM, KNN, MNL, DT, BM DNN
17 Ermagun et al. (2015) [28] MC $4,700$ 2 - NL, RF RF
18 Jahangiri et al. (2015) [29] MC N.A. 6 - KNN, SVM, DT, BAGGING, RF, MNL RF
19 Tang et al. (2015) [30] MC $72,536$ 2 - DT, MNL DT
20 Omrani (2015) [31] MC $9,500$ 4 - DNN, RBFNN, MNL, SVM DNN
21 Shafique et al. (2015) [32] MC $1,968$ 4 - SVM, BOOSTING, DT, RF RF
21 Shafique et al. (2015) [32] MC $1,488$ 4 - SVM, BOOSTING, DT, RF RF
21 Shafique et al. (2015) [32] MC $2,754$ 4 - SVM, BOOSTING, DT, RF RF
22 Shukla et al. (2015) [33] MC $100,000$ 2 - DNN, DT DNN
23 Sekhar and Madhu (2016) [34] MC $4,976$ 3 - RF, DT, MNL RF
24 Hagenauer and Helbich (2017) [5] MC $230,608$ 8 - MNL, DNN, NB, SVM, CTs, BOOSTING, BAGGING, RF RF
25 Paredes et al. (2017) [35] CO $15,211$ 5 - MNL, BOOSTING, DT, SVM, RF RF
26 Hillel et al. (2018) [36] MC N.A. 8 - DNN, BAGGING, BOOSTING, KNN, GLM, BM, RF, SVM BOOSTING
27 Golshani et al. (2018) [37] MC $9,450$ 2 - MNL, DNN DNN
@@ 24 Hagenauer and Helbich (2017) [5] MC $230,608$ 8 - MNL, DNN, NB, SVM, CTs, BOOSTING, BAGGING, RF RF
29 Wang and Ross (2018) [6] MC $51,910$ 2 - BOOSTING, MNL BOOSTING
@@
31 Cheng et al. (2019) [7] MC $7,276$ 4 - RF, SVM, BOOSTING, MNL RF
32 Zhou et al. (2019) [38] MC $30,000$ 8 - MNL, KNN, DT, SVM, BM, BOOSTING, BAGGING, RF BAGGING
33 Wang et al. (2020) [9] MC $8,418$ 8 - MNL, NL, DNN, SVM, BM, KNN, BOOSTING, GLM DNN
33 Wang et al. (2020) [9] MC $,92$ 929 8 - MNL, NL, DNN, SVM, BM, KNN, BOOSTING, GLM DNN
34 Wang et al. (2020) [8] MC $80,000$ 2 - MNL, DNN DNN
34 Wang et al. (2020) [8] MC $8,418$ 2 - MNL, DNN DNN
35 Wang et al. (2020) [39] MC $8,418$ 3 - NL, MNL, DNN DNN
@@ 4.1 Treatment dimension: models and model families
Table LABEL:table:list_classifiers summarizes the list of 105 individual models (Column 1) from 12 model families (Column 2). Instead of analyzing single models, we create a two-tier structure for models and model families, thus recognizing the richness within each model family. For example, the DNN family includes 16 DNNs with different architectures, thus enabling us to analyze the variation across the DNN architectures. While far from exhaustive, Table LABEL:table:list_classifiers has included the relatively representative models and the most relevant ones for travel behavioral analysis. For example, DCMs are highly relevant given their consistent use in the travel demand analysis, and DNNs are also highly relevant due to their rising popularity in many subdomains in transportation [40, 9, 39].
@@ 2. Deep Neural Networks (15 Models)
avNNet_R DNN Neural network with random seeds with averaged scores; [41] R Caret avNNet
@@
monmlp_R DNN MLP with monotone constraints [42] R Caret monmlp
@@ 3. Discriminant Analysis (12 Models)
pda_R DA Penalized discriminant analysis (PDA) with shrinkage penalty coefficients [43] R mda gen.ridge
mda_R DA Mixture discriminant analysis (MDA) where the number subclass is tuned to 3 [44] R mda
rda_R DA Regularized discriminant analysis (RDA) with regularized group covariance matrices [45] R klaR
hdda_R DA High dimensional discriminant analysis (hdda) assuming each class in a Gaussian subspace [46] R HD
@@ 4. Bayesian Models (7 Models)
BayesNet_W BM Bayes network models by hill climbing algorithm [47] Weka BayesNet
@@ 6. K Nearest Neighbors (4 Models)
lBk_1_W KNN KNN classifier with number of neighbors equal to 1 (brute force searching and Euclidean distance) [48] Weka lBk
lBk_5_W KNN KNN classifier with number of neighbors equal to 5 (brute force searching and Euclidean distance) [48] Weka lBk
@@ 7. Decision Tree (14 Models)
C5.0Rules_R DT Rule-based models using Quinlan’s C5.0 algorithm [49] R Caret C5.0Rules
ctree_R DT Conditional inference trees [50] R Caret ctree
@@
ExtraTree_P DT Tree classifier with best splits and features chosen from random splits and randomly selected features [51] Python sklearn ExtraTreeClassifier
@@
HoeffdingTree_W DT An incremental tree with inductive algorithm. [52] Weka HoeffdingTree
@@
DecisionTable_W DT Simple decision table majority classier that uses BestFirst as search method [53] Weka DecisionTable
@@ 8. Generalized Linear Models (10 Models)
Logistic_W GLM Logistic regression model with a ridge estimator [54] Weka Logistic
SimpleLogistic_W GLM Linear logistic regression models fitted by using LogitBoost [55] Weka SimpleLogistic
@@
Passive Aggressive_P GLM Passive-aggressive algorithms for classification with hinge loss [56] Python sklearn PassiveAggressiveClassifier
@@ 12. Boosting (23 Models)
AdaBoost_P BOOSTING AdaBoost classifier. The DecisionTree with maximum depth =10 is set as the base estimator. [57] Python sklearn AdaBoostClassifier
@@
Gradient Boosting_P BOOSTING An additive model trained in a forward stage-wise fashion [58] Python sklearn GradientBoostingClassifier
@@ 4.2 Contextual factors
The contextual variables are data sets, sample sizes, and choice categories, which can also influence the model performance besides the model choice. The data sets include NHTS2017, LTDS2015, and SGP2017. NHTS2017 refers to the national household travel survey 2017, collected by the Federal Highway Administration [^3] Available at https://nhts.ornl.gov/, which provides travel, vehicle, and individual information across the United States. LTDS2015 refers to the London travel demand survey 2015, which was collected in London with the history of trips from April 2012 to March 2015 and augmented with mode-specific level-of-service variables (e.g. in-vehicle travel time, public transport fares, fuel cost, etc.) [59]. SGP2017 was collected in 2017 in Singapore through a stated preference survey for travel mode choice [60, 61, 62], in which the experiments followed the orthogonal survey design [63]. SGP2017 contains information about travel mode choices, socio-demographics, and alternative specific variables. The three data sets are chosen to cover a variety of data collection procedures, geographical locations, and local contexts. The NHTS2017 was collected through revealed preference surveys by the US government, LTDS2015 by combining a transit agency’s survey with simulation-based travel information, and SGP2017 by a standard stated preference survey. The three datasets cover the continents of America, Europe, and Asia, thus jointly creating a geographically diverse set.
@@
The choice categories include three travel behavior variables: travel modes, trip purposes, and car ownership, which are amongst the most common choice categories from the literature. Travel mode choice is available for all three data sets, while trip purposes and car ownership are only available for the NHTS2017 data. In the NHTS2017 data, the initial 21 travel mode categories are aggregated into six major modes: walk and bike, car, SUV, van and track, public transit, and others. The NHTS2017 data has five trip purposes, including home-based work, home-based shopping, home-based social, home-based others, and none-home-based trips. The car ownership in the NHTS2017 data also has five categories, ranging from zero to more than three cars. The outputs from the three data sets are summarized in Figure 1. In the NHTS2017 data, the most common travel mode is car, accounting for 45.3%, while only 1.8% of the trips are taken by public transit. In the LTDS2015 and SGP2017 data sets, the largest mode share is driving, which accounts for 44.1% and 44.7%, respectively. Public transit is the second largest mode share in London and Singapore, accounting for 35.3% and 23.0% of the total trips.
@@
Figure 1: Output distributions in three datasets
@@ 5.2 Examining performance across model families and individual models
Figures 2 and 3 present visualizations of prediction accuracy for 12 model families and 105 individual models. Both figures are sorted based on average prediction accuracy, from highest to lowest. In Figure 2, the x-axis represents model families, while the y-axis represents prediction accuracy. Each white dot represents an experiment’s accuracy, and the grey areas indicate density distributions of predictive performance. The blue, red, and green dots represent the mean, median, and maximum prediction accuracy, respectively. Figure 3 extends the 12 model families to 105 classifiers using a similar format, replacing the violin plots with simple blue bars to depict performance variance. In this figure, the blue and red dots respectively represent the mean and median prediction accuracy. Additionally, Appendix IV contains an analysis of literature data, yielding similar findings to the experiment data presented here.
Figure 2: Prediction accuracy of 12 model families
@@
Figure 2 illustrates the relative prediction accuracy of different model families. Among them, ensemble methods and DNNs exhibit relatively high prediction accuracy, while the DCM family ranks 7th out of 12. The DNN family, comprising 15 classifiers, achieves mean and median accuracies of 57.79% and 56.77% respectively. The three ensemble methods, namely random forests, boosting, and bagging, also demonstrate high accuracy with mean values of 57.05%, 57.03%, and 55.42% respectively. Despite incorporating more complex error structures in the NL and MXL models, the DCMs still fail to outperform the ensemble methods and DNNs. The superior prediction accuracy of ML models can be attributed to their strong approximation power [64] and effective regularization techniques [65, 66]. This visualization further confirms the findings from our previous tournament model, providing additional ranking information for each model family.
Figure 3: Prediction accuracy of 105 models
@@
Figure 2 also reveals significant variation and even multi-modal distributions in predictive performance. For instance, the DNN family exhibits an average accuracy of approximately 57.79%, while its maximum prediction accuracy reaches around 74.61%. Similarly, the average accuracy of the DCM family stands at 54.43%, which is 15.37% lower than its maximum value of 69.8%. Remarkably, each model family can achieve maximum values in the range of 70-80%, surpassing the average by 20-30 percentage points. Several model families display explicit multi-modal distributions, including DCM and BAGGING. The wide variation and multi-modal distributions in prediction accuracy visually demonstrate that the choice of models may not be the most crucial factor in determining the final predictive performance. This observation aligns with our previous finding regarding the substantial uncertainty inherent in such model comparisons.
@@
Figure 3 demonstrates that individual models do not deviate significantly from the average performance of their respective model families, and substantial variation persists within each of the 105 individual models. Among these models, the top performers are LogitBoosting and GradientBoosting from the BOOSTING family, followed by DNNs such as nnet, avNNet, and monmlp, which rank 4th, 5th, and 10th out of 105 models. Within the DCM model family, MNL, NL, and MXL (highlighted in red) exhibit similar performance at the medium to lower end of the spectrum among the 105 models. Their average prediction accuracies are 54.50%, 55.01%, and 52.58%, respectively, only slightly deviating from the DCM family’s average. However, each individual model represents a significant variation, with performance varying by at least 20 percentage points, as indicated by the long blue bars. The fact that MXL performs worst may come as a surprise to readers, given that in estimation, the inclusion of random heterogeneity almost invariably leads to big gains in performance. However, here, we are talking about prediction performance, which implies averaging again over the random components, and this removes much of the differences between MXL and non-mixture DCM models.